A teacher selects 3 students from the 22 students in a physics class to represent Your school at the science fair. In how many different ways can the teacher choose?
Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(22,3) = (22!)/(3!*(22-3)!)

C(22,3) = (22!)/(3!*19!)

C(22,3) = (22*21*20*19!)/(3!*19!)

C(22,3) = (22*21*20)/(3!)

C(22,3) = (22*21*20)/(3*2*1)

C(22,3) = (9240)/(6)

C(22,3) = 1540

So there are 1540 different ways.